Fermi National Laboratory
 

Accelerator Update for July 23, 2007 - July 25, 2007

 

Monday July 23

The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (366.1E10), with the Antiproton Source (Pbar) stacking (32.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams, and with MiniBooNE and the Meson MTest experiments taking beam. NuMI will remain off until after the summer shutdown.

At 9:15 AM, during the first transfer of antiprotons from Pbar to the Recycler, the P1 transfer line tripped off due to a quadrupole power supply (Q713) not following its reference. Everything reset okay, but then the P1 line tripped off on all the rest of the transfers due to Q713. Operations contacted an expert who made a small change to a timing card.

At 9:47 AM, the TeV suffered a sector BA and BB quench. Operators and TeV experts investigated.

Operations reestablished beam to MTest at 10:28 AM.

A separator expert reported at 10:43 AM that he would be replacing a separator power supply at TeV sector B0.

MI and Pelletron experts began conducting studies at 10:51 AM.

Operations prepared CDF and D0 for access at 11:16 AM.

CDF came out of access at 1:05 PM.

D0 came out of access at 1:52 PM.

Operations began the TeV turn on at 2:07 PM.

Operations ramped the TeV at 3:18 PM.

Operations put the TeV through a dry squeeze at 3:31 PM.

Operations reported at 3:57 PM that the Recycler had stashed antiprotons three times during the day shift.

MTest stopped taking beam for the night at 6:03 PM.

Operations put the TeV into shot setup at 9:37 PM.

At 10:54 PM, Operations reported that while loading antiprotons from MI to the TeV, the P1 Line was again tripping off. Operations contacted experts.

Operations established store 5573 at 11:44 PM with an initial luminosity of 110.68E30. Operators reported several problems during this setup; beam positions were bad and some of the collimators didn’t make it to their correct positions. Operators had to scrape the beam twice and retune.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Tuesday July 24

The midnight shift began with Operations monitoring store 5573 (94.95E30) and the stash (2.8E10), with Pbar stacking (81.5E10), and with MiniBooNE taking beam.

A Booster expert began a 40-minute study at 1:41 AM.

At 2:10 AM, a Linac RF station (LRF3) tripped off. It would not reset remotely. It took operators several tries to reset locally.

At 3:25 AM, Operations began stashing antiprotons. The third transfer would not extract from the Accumulator. Operators put Pbar back into stacking mode and tried to cool the beam, but lost nearly 20mA of antiprotons. Operations contacted an expert.

Operations prepared the Recycler to stash again at 5:44 AM, but a RF station (ARF4) did not produce any voltage. Operators put Pbar back into stacking mode and called in an expert.

Operations and an expert prepared the Recycler to stash at 7:01 AM. Five transfers were completed.

MTest resumed taking beam at 7:36 AM.

Operations reported at 8:01 AM that the Recycler had stashed twice during the midnight shift.

At 9:05 AM, BRF13 tripped off on a ground fault indication. It reset, but then tripped off again. Operators bypassed the station and lowered the intensity. Expert investigated.

Operations reported at 9:59 AM that Preaccelerator experts began switching from the H- Source to the I- Source.

TeV experts began an end-of-store study at 10:02 AM.

At 11:01 AM, an RF expert reported that BRF13 had a Power Amplifier LCW leak.

At 11:57 AM, LRF3 tripped off. Operators and experts investigated and discovered that someone was using a two-way radio near the station.

The TeV end-of-store study ended at 11:58 AM.

Operations prepared Booster for an access at 12:05 PM.

Operations terminated store 5573 at 12:08 PM.

Operations prepared D0 for an access at 12:17 PM.

A Linac expert reported at 12:33 PM that he had brought LRF3 back online.

RF experts reported at 12:56 PM that they were swapping out the bias power supply for MIRF16.

Recycler and Linac experts began separate studies at 1:50 PM.

D0 came out of access at 2:02 PM.

The Linac studies ended at 2:19 PM.

At 2:46 PM, Operations reported a Pbar readback problem. The Pbar consoles were not communicating. Operators contacted controls experts.

Booster came out of access at 3:10 PM.

Controls experts reported at 3:40 PM that they found a switch at AP-10 that had an uplink failure. They replaced the switch.

Operations reported at 3:52 PM that Booster was up and running and they had established beam to Pbar, MiniBooNE, and MTest.

Operations reported at 4:03 PM that the Recycler had stashed once during the day shift.

Operations reported at 4:46 PM that the BRF13 LCW leak had been repaired, but the station needed more time to dry out before it could be run.

Operations disabled the MTest beam at 6:03 PM.

TeV studies ended. Operations put the TeV into shot setup at 6:20 PM.

At 9:13 PM, LRF3 tripped off. It reset.

Operations established store 5577 at 9:39 PM with an initial luminosity of 147.93E30.

At 10:10 PM, LRF3 tripped off. It reset.

A Recycler expert took some damper system measurements at 11:18 PM.

Operations reported at midnight that the Recycler had stashed three times during the midnight shift.

Wednesday July 25

The midnight shift began with Operations monitoring store 5577 (95.46E30) and the stash (76.8E10), with Pbar stacking (17.1E10), and with MiniBooNE taking beam.

At 2:03 AM, Pbar tripped off. Operators investigated and discovered a correction magnet power supply (HTF17) had tripped off. When operators turned it on it ramped up to 10 amps, which was much higher that it normally ran, and this caused the bad beam position and trip. Apparently, a program used to control the beam didn’t know the supply was off and kept increasing HTF17’s current up to its maximum trying to move the beam. The trip was reset at 3:07 AM.

MTest resumed taking beam at 6:34 AM.

Operations reported at 7:57 AM that the Recycler had stashed twice during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Recycler, MI, and Pbar experts to conduct studies today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/25/2007    email Fermilab